OpenStudy (anonymous):

1. Ibuprofen, C13H18O2, is the active ingredient in many non-prescription pain relievers. (a) If the tablets in a bottle contain a total of 13 g of ibuprofen, how many moles of ibuprofen are in the bottle? Work N= mass/mr N=13/206.29g N=0.0630n Answer 0.0630n (b) How many atoms of ibuprofen are in the bottle?

OpenStudy (aaronq):

so you got the moles to go from moles to molecules divide them moles by avogadros number \(molecules=\dfrac{moles}{N_A}\)

OpenStudy (aaronq):

if it wants ATOMS and not just molecules, then you multiply the number of molecules by the number of atoms 1 molecules of ibuprofen consists of.

OpenStudy (anonymous):

so which ones am i divinging? all of the percentages by the moles and adding them?!

OpenStudy (aaronq):

moles/avogadros number

OpenStudy (anonymous):

i dont know what that means!

OpenStudy (aaronq):

but are they asking for molecules of ibuprofen or total number of atoms?

OpenStudy (anonymous):

total number of atoms in the entire bottle of ibuprofen

OpenStudy (aaronq):

N= mass/mr N=13/206.29g N=0.0630n Answer 0.0630n this is moles so moles/avogadros number = \(\dfrac{0.0630}{6.022*10^{22}}\)

OpenStudy (aaronq):

i meant to the power of 23, as in "\(6.022*10^{23}\)"

OpenStudy (anonymous):

thank you!